### v1

@@ Abstract
Agentic systems offer a potential path to solve complex clinical tasks through collaboration among specialized agents, augmented by tool use and external knowledge bases. Nevertheless, for chest X-ray (CXR) interpretation, prevailing methods remain limited: (i) reasoning is frequently neither clinically interpretable nor aligned with guidelines, reflecting mere aggregation of tool outputs; (ii) multimodal evidence is insufficiently fused, yielding text-only rationales that are not visually grounded; and (iii) systems rarely detect or resolve cross-tool inconsistencies and provide no principled verification mechanisms. To bridge the above gaps, we present RadAgents, a multi-agent framework for CXR interpretation that couples clinical priors with task-aware multimodal reasoning. In addition, we integrate grounding and multimodal retrieval-augmentation to verify and resolve context conflicts, resulting in outputs that are more reliable, transparent, and consistent with clinical practice.
@@
We use the following public datasets: MIMIC-CXR-JPG (Johnson et al., 2019), MS-CXR (Boecking et al., 2022), and MS-CXR-T (Bannur et al., 2023), which are accessible under their respective data use agreements. We plan to release the code after obtaining organizational approval.
@@ 1 Introduction
Chest X-ray (CXR) imaging is a cornerstone of pulmonary screening, diagnosis, and follow-up, accounting for the largest share of diagnostic radiology examinations performed worldwide each year (Cid et al., 2024). Yet systematic assessment of thoracic structures remains labor-intensive, imposing a substantial time burden on radiologists (Fallahpour et al., 2025). The gradual introduction of AI into clinical practice shows promise for alleviating this workload (Zhang et al., 2024; Tanno et al., 2025). However, prevailing systems fall short on complex multimodal reasoning, such as integrating findings across disparate image regions, which is central to radiologists’ practice; most adhere to end-to-end designs in which the visual encoder executes a single, front-end pass and subsequent reasoning proceeds purely in text (Wang et al., 2025). This encode-once, text-only paradigm decouples the reasoning trajectory from evolving visual evidence, leading to failures on tasks that require iterative re-inspection, precise measurements, and cross-comparisons (Liu et al., 2025).
A promising path for clinical reasoning is to augment large language models, including multimodal variants, with external tools. By delegating perceptual and classification subtasks such as organ or region segmentation and disease classification to validated modules, the language model can focus on planning and synthesis. Several agentic frameworks already explore this idea, from fine tuned tool use (Li et al., 2024; Nath et al., 2025) to pipelines that call commercial models such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), although multiple agent coordination and/or debate introduces considerable computational overhead. In CXR interpretation, RadFabric (Chen et al., 2025) integrates diagnostic agents with a separate reasoning agent, and MedRAX (Fallahpour et al., 2025) expands task coverage by incorporating additional task specific models. Despite gains over single-model baselines, integration and reasoning steps are often opaque and not aligned with clinical workflow, which undermines trust and creates safety risks, and most systems still lack explicit verification and conflict resolution.
In this paper, we introduce RadAgents, a training free framework (that is, no weight updates), designed to be trustworthy and extensible for CXR interpretation, especially in cases that require complex multimodal reasoning. Evaluated on three datasets across two core tasks: report generation and visual question answering (VQA), RadAgents achieves state of the art performance, surpassing the strongest baselines by 10.2% on MS-CXR, 29.6% on MIMIC-CXR, and 21.5% on MS-CXR-T, underscoring the value of incorporating clinically-aware workflows. Our agentic design, enhanced with multimodal retrieval-augmentation, delivers consistent gains: an average improvement of 8.0% over the no-retrieval variant, demonstrating its effectiveness in mitigating context conflicts and producing more reliable results.
@@ 2 Methods
RadAgents is a multi agent system with seven specialized agents (Figure 1). Five implement the clinical ABCDE review scheme (Hodler et al., 2019): Airway, Breathing, Circulation, Diaphragm, and Everything else. In addition, an Orchestrator agent analyzes each query and routes tasks to the appropriate specialists with the required patient context (for example, imaging view and prior studies), and a Synthesizer agent integrates their outputs, resolves conflicts, and produces the final output.
This design confines context to task specific compartments, reducing the information each agent must process and simplifying context compression for synthesis. It also allows parallel execution, lowering latency for long reasoning. For common and clinically significant CXR findings such as cardiomegaly and pleural effusion, we curate radiologist like workflows that guide tool selection and clinically grounded reasoning (see the demonstration example in Appendix A). Beyond these predefined templates, the language model handles queries that fall outside fixed workflows, plans tool use, and enables flexibility in interpretation.
@@ 2.1 Task-aware Sub-Agents
Each sub-agent, also called the ABCDE agent, has a defined purpose and domain of expertise. Each is governed by a custom system prompt derived from clinical guidelines and maintains its own context window. The main scope and objectives of them are:
@@ 2.3 Global Controller Module
Query analysis. Given a query, the Orchestrator drafts a high level plan, selects the relevant subagents, and chooses the reasoning mode: ReAct when no workflow is specified (Yao et al., 2023), or Plan-and-Execute (P&E) (WANG et al., 2023) when a workflow template is available. This keeps the system language driven and adaptable across queries.
Tool cards. We employ a suite of models as tools for distinct CXR tasks: CheXagent (Chen et al., 2024b) for VQA, MAIRA-2 (Bannur et al., 2024) for grounding, the CheXpert Plus report generator (Chambon et al., 2024), and classification and organ segmentation models from TorchXRayVision (Cohen et al., 2022). In addition, we include unique programming tools that return zoomed-in quarter patches or serve for measurement and calculation purposes.
Context verifier for tools. No tool is perfect, as their capabilities are constrained by model size and training data. When uncertainty arises, we trigger a verification step in which an advanced multimodal LLM serves as a judge (Chen et al., 2024a), filtering out incorrect outputs such as erroneous masks.
Retrieval-augmented conflic resolution. Tool outputs can conflict. On the Synthesizer side, we apply Visual Retrieval-Augmented Generation (V-RAG) (Chu et al., 2025): the agent retrieves clinically similar chest radiographs (based on image embeddings from Rad-DINO (Perez-Garcia et al., 2025)) and accompanying context like patient notes and uses these exemplars to adjudicate discrepancies among tools (Figure 2). This mirrors routine radiologic practice, in which clinicians consult similar cases and content to calibrate interpretation.
Figure 2: Resolving the conflicts via V-RAG.
@@ 3.1 Experimental Setup
Baselines. We instantiate all agents with GPT-4o. For comparison, we include (1) GPT-4o (Hurst et al., 2024), (2) GPT-4o with ReAct, where tools are available via function calling but without explicit workflow steering, and (3) two medical specialist models, CheXagent and MedGemma (Sellergren et al., 2025). Unless otherwise noted, the number of retrieved exemplars for V-RAG is set to $k=3$ (see Appendix F for an ablation on $k$). We report results for two variants of RadAgents, with and without V-RAG.
Metrics. For VQA (E&A) and report generation, we use standard CXR text metrics (explanation in Appendix B): RadGraph F1 (Jain et al., 2021), CheXbert macro F1 across 14 labels (Smit et al., 2020), RaTE (Zhao et al., 2024), and GREEN (Ostmeier et al., 2024). Because the outputs are sentences, these metrics capture clinical correctness and consistency. For VQA (C&P), where the output is one of three choices, we report accuracy.
Table 1: Evaluation on VQA (E&A) and report generation with baselines and RadAgents.
@@ 3.2 Existence and Attributes
The VQA questions cover seven common findings in CXR: atelectasis, cardiomegaly, consolidation, edema, lung opacity, pleural effusion, and pneumothorax, derived from the standard test split of MS–CXR. Each image receives the following prompt: {infobox}<image> Describe if [finding] is present; if present, describe [attributes].
@@ 3.3 Comparison and Progression
We use MS–CXR–T to assess stability, improvement, or worsening of a specific positive finding (consolidation, edema, pleural effusion, or pneumothorax). We only retain cases where the metadata indicates a consensus among human reviewers. We pose a comparative question that explicitly references the prior study. The prompt template is: {infobox}Given current image <image>, and previous image <image>, decide if [finding] is improving, stable, or worsening.
@@ 3.4 Report Generation
We construct a MIMIC–CXR subset aligned with MS–CXR identities so that findings queried in VQA are represented in the corresponding reports. Prompts request generation of the Findings section, and all agents are activated by default. The prompt combines the template from Section 3.2 with a curated list of clinically significant findings, following prior work (Tu et al., 2024; Peng et al., 2025); details appear in Appendix D.
@@
Figure 3: VQA results regrading progression.
@@ 4 Discussion
In this work, we propose the first radiologist-like agentic system, which demonstrates superior performance in CXR interpretation. Further improvements are required for broader application: (1) the context verifier can only detect errors but cannot modify outputs, especially for visual evidence such as bounding boxes, so stronger supporting models are needed; (2) current experiments are limited to frontal-view images, as most existing tools are incompatible with lateral views; (3) prompting strategies could be further optimized; and (4) the framework could be extended to additional modalities, such as CT and MRI.
@@ Appendix B Clinical Evaluation Metrics
Evaluating the quality of generated radiology reports is non-trivial. Early works adopted general-domain natural language processing metrics such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002). While these metrics are widely used for text evaluation, they treat differences in wording the same as clinically significant errors, failing to reflect medical accuracy. To address this limitation, clinically informed evaluation metrics, such as CheXbert (Smit et al., 2020), RadGraph (Jain et al., 2021), GREEN (Ostmeier et al., 2024), and RaTEScore (Zhao et al., 2024), have been proposed to better assess clinical correctness and utility. CheXbert is based on multi-label classification results for 5 or 13 diseases (along with one extra “normal” label). RadGraph considers literal entity agreement considering the positive or negative context of each entity. GREEN judges recall and precision errors by LLM prompting. RaTEScore is inspired by RadGraph but less sensitive to phrasing by an F1-like computation which allows semantic matching between entities based on a cosine similarity. The metrics are computed using their official and standardized implementations: RadGraph-F1111https://pypi.org/project/radgraph/0.1.2/, CheXbert-F1222https://pypi.org/project/f1chexbert/, RaTE Score [^3] https://pypi.org/project/RaTEScore/0.5.0/, and GREEN [^4] https://pypi.org/project/green-score/0.0.8/.
@@ Appendix E Reasoning modes and agentic workflows
We first introduce clinically aware interleaved reasoning, which identifies significant findings through context assembly from metadata, targeted tool use, and self guided reasoning. We then outline the RadAgents workflow, driven by prompt based task decomposition, strategic tool selection, and report synthesis.
@@ Appendix F Details of V-RAG
We retrieve images and their textual descriptions that align with the features of target medical images following (Chu et al., 2025). These references, rich in visual and textual medical details, guide response generation. To obtain embeddings, we use Rad-DINO, which provides robust representations across diverse CXR image types. For each image $X_{img}$, we extract its embedding $E_{img}=\mathbf{R}^{d}$, with $d=768$, and store them in the embedding memory $\mathcal{M}$.
For efficient retrieval during inference, we build $\mathcal{M}$ using FAISS  [^5] https://github.com/facebookresearch/faiss, a GPU-accelerated vector search system. We employ approximate kNN with the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2018), enabling retrieval of the top-$k$ most similar images in $\mathcal{M}$.
@@
We study how retrieval quality changes with the number of retrieved studies $k$ with the sampled 100 cases from the MS-CXR test set used in this study. For each setting, we compute two metrics: helpful-rate, the percentage of retrieved studies that improve the answer, and harmful-rate, the percentage that hurt the answer. As shown in Table 3, the helpful-rate increases from 0.48 at $k=1$ to 0.65 at $k=5$, while the harmful-rate also rises from 0.09 to 0.20. This illustrates a trade-off: retrieving more studies provides greater chances of including helpful evidence but also increases the risk of introducing misleading content. To balance these effects, we choose $k=3$ by default, achieving a helpful-rate of 0.62 with a moderate harmful-rate of 0.14.

### v2

@@ Abstract
Agentic systems offer a potential path to solve complex clinical tasks through collaboration among specialized agents, augmented by tool use and external knowledge bases. Nevertheless, for chest X-ray (CXR) interpretation, prevailing methods remain limited: (i) reasoning is frequently neither clinically interpretable nor aligned with guidelines, reflecting mere aggregation of tool outputs; (ii) multimodal evidence is insufficiently fused, yielding text-only rationales that are not visually grounded; and (iii) systems rarely detect or resolve cross-tool inconsistencies and provide no principled verification mechanisms. To bridge the above gaps, we present RadAgents, a multi-agent framework that couples clinical priors with task-aware multimodal reasoning and encodes a radiologist-style workflow into a modular, auditable pipeline. In addition, we integrate grounding and multimodal retrieval-augmentation to verify and resolve context conflicts, resulting in outputs that are more reliable, transparent, and consistent with clinical practice.
@@
We use the following public datasets: MIMIC-CXR-JPG (johnson2019mimic), MS-CXR (boecking2022making), and MS-CXR-T (bannur2023learning), which are accessible under their respective data use agreements. We plan to release the code after obtaining organizational approval.
@@ 1 Introduction
Chest X-ray (CXR) imaging is a cornerstone of pulmonary screening, diagnosis, and follow-up, accounting for the largest share of diagnostic radiology examinations performed worldwide each year (cid2024development). Yet systematic assessment of thoracic structures remains labor-intensive, imposing a substantial time burden on radiologists (fallahpourmedrax). The gradual introduction of AI into clinical practice shows promise for alleviating this workload (zhang2024generalist; tanno2025collaboration). However, prevailing systems fall short on complex multimodal reasoning, such as integrating findings across disparate image regions, which is central to radiologists’ practice; most adhere to end-to-end designs in which the visual encoder executes a single, front-end pass and subsequent reasoning proceeds purely in text (wang2025multimodal). This encode-once, text-only paradigm decouples the reasoning trajectory from evolving visual evidence, leading to failures on tasks that require iterative re-inspection, precise measurements, and cross-comparisons (liu2025more).
A promising path for clinical reasoning is to augment large language models, including multimodal variants, with external tools (lu2025octotools). By delegating perceptual and classification subtasks such as organ or region segmentation and disease classification to validated modules, the language model can focus on planning and synthesis. Several agentic frameworks already explore this idea, ranging from training small models for limited tool use (li2024mmedagent; nath2025vila) to pipeline systems that invoke general-purpose models for more flexible operations (jiang2025medagentbench; schmidgall2024agentclinic), although multiple agent coordination and/or debate introduces considerable computational overhead. In CXR interpretation, RadFabric (chen2025radfabric) integrates diagnostic agents with a separate reasoning agent, and MedRAX (fallahpourmedrax) expands task coverage by incorporating additional task specific models. Despite gains over single-model baselines, integration and reasoning steps are often opaque and not aligned with clinical workflow, which undermines trust and creates safety risks, and they still lack explicit verification and conflict resolution.
To address these limitations, we introduce RadAgents, a framework designed for complex multimodal reasoning in CXR, which has three primary technical contributions:
  • We introduce clinically aware interleaved reasoning, a multimodal loop that combines clinical priors and study metadata with targeted image operations (segment, measure, localize, compare), so each hypothesis triggers tool use and returns inspectable evidence, mirroring how radiologists proceed from observation to measurement, comparison with prior studies, and conclusion (Fig. 6 and Appendix E).
  • We propose a training-free multi-agent design, where five sub-agents operate in clean, task-scoped context compartments, coordinated by an Orchestrator and reconciled by a Synthesizer. This preserves visual grounding, supports parallel execution, and composes measurement, localization, characterization, comparison, and diagnosis (Figure 1).
  • We add a lightweight context verifier and visual retrieval-augmented generation to detect, surface, and resolve inconsistencies before reporting, yielding more clinically aligned outputs.
  • RadAgents achieves state-of-the-art performance, outperforming strong baselines by 10.2% on MS-CXR, 29.6% on MIMIC-CXR, and 21.5% on MS-CXR-T. Multimodal retrieval further contributes an average 8.0% boost over the no-retrieval variant, mitigating context conflicts and improving reliability.
@@ 2 Methods
RadAgents is a multi agent system with seven specialized agents (Figure 1). Five implement the clinical ABCDE review scheme (hodler2019diseases): Airway, Breathing, Circulation, Diaphragm, and Everything else. In addition, an Orchestrator agent analyzes each query and routes tasks to the appropriate specialists with the required patient context (for example, imaging view and prior studies), and a Synthesizer agent integrates their outputs, resolves conflicts, and produces the final output.
This design confines context to task specific compartments, reducing the information each agent must process and simplifying context compression by having each sub-agent produce an initial summary for downstream synthesis. It also allows parallel execution, lowering latency for long reasoning. For common and clinically significant CXR findings such as cardiomegaly and pleural effusion, we curate radiologist-like workflows, the predefined templates within RadAgents, to guide tool selection and clinically grounded reasoning (see the demonstration example in Appendix A). For out-of-template queries, the system invokes workflow-free reasoning, preserving flexibility. The design is extensible: new templates (e.g., reasoning or tool-chains) can be added, and some can generalize to tasks of similar scope or category.
@@ 2.1 Task-aware Sub-Agents
Each sub-agent, also called the ABCDE agent, has a defined purpose and domain of expertise. Each is governed by a custom system prompt derived from clinical guidelines and maintains its own context window (See details ). The main scope and objectives of them are:
@@ 2.3 Global Controller Module
Query analysis. Given a query, the Orchestrator drafts a high level plan, selects the relevant subagents, and chooses the reasoning mode: ReAct when no workflow is specified (yao2023react), or Plan-and-Execute (P&E) (wang2023lim) when a workflow template is available. This keeps the system language driven and adaptable across queries.
Tools. We employ a suite of models as tools for distinct CXR tasks: CheXagent (chen2024chexagent) for VQA, MAIRA-2 (bannur2024maira) for grounding, the CheXpert Plus report generator (chambon2024chexpert), and classification and organ segmentation models from TorchXRayVision (cohen2022torchxrayvision). In addition, we include unique programming tools that return zoomed-in quarter patches or serve for measurement and calculation purposes.
Context verifier. No tool is perfect, as their capabilities are constrained by model size and training data. When uncertainty arises, we trigger a verification step in which an advanced multimodal LLM serves as a judge (chen2024mllm), filtering out incorrect outputs such as erroneous masks.
Retrieval-augmented conflict resolution. Tool outputs can conflict. On the Synthesizer side, we apply Visual Retrieval-Augmented Generation (V-RAG) (chu2025reducing): the agent retrieves clinically similar chest radiographs (based on image embeddings from Rad-DINO (perez2025exploring)) and accompanying context like patient notes and uses these exemplars to adjudicate discrepancies among tools (Figure 2 and Appendix F). This mirrors routine radiologic practice, in which clinicians consult similar cases and content to calibrate interpretation.
Figure 2: Resolving the conflicts via V-RAG (chu2025reducing).
@@ 3.1 Experimental Setup
Baselines. We instantiate all agents with GPT-4o. For comparison, we include (1) GPT-4o (hurst2024gpt), (2) GPT-4o with ReAct, where tools are available via function calling but without explicit workflow steering, (3) GPT-4o with monolithic Workflow where a MLLM executing ABCDE analysis end-to-end with the same tools, retrieval, and budges as RadAgents, and (4) two medical specialist models, CheXagent and MedGemma (sellergren2025medgemma). Unless otherwise noted, the number of retrieved exemplars for V-RAG is set to $k=3$ (see Appendix F for an ablation on $k$). We report results for two variants of RadAgents, with and without V-RAG.
Metrics. For VQA (E&A) and report generation, we use standard CXR text metrics (explanation in Appendix B): RadGraph F1 (jain1radgraph), CheXbert macro F1 across 14 labels (smit2020combining), RaTE (zhao2024ratescore), and GREEN (ostmeier2024green). Because the outputs are sentences, these metrics capture clinical correctness and consistency. For VQA (C&P), where the output is one of three choices, we report accuracy.
Table 1: Evaluation on VQA (E&A) and report generation with baselines and RadAgents. The four metrics are commensurate (each normalized to the range $[0,1]$).
@@ 3.2 Existence and Attributes
The VQA questions cover seven common findings in CXR: atelectasis, cardiomegaly, consolidation, edema, lung opacity, pleural effusion, and pneumothorax, derived from the standard test split of MS–CXR. Each image receives the following prompt (details are stated in the Appendix D):
{infobox}
<image> Describe if [finding] is present; if present, describe [attributes].
@@ 3.3 Comparison and Progression
We use MS–CXR–T to assess stability, improvement, or worsening of a specific positive finding (consolidation, edema, pleural effusion, or pneumothorax). We only retain cases where the metadata indicates a consensus among human reviewers. We pose a comparative question that explicitly references the prior study. The prompt template is:
{infobox}
Given current image <image>, and previous image <image>, decide if [finding] is improving, stable, or worsening.
@@ 3.4 Report Generation
We construct a MIMIC–CXR subset aligned with MS–CXR identities so that findings queried in VQA are represented in the corresponding reports. Prompts request generation of the Findings section, and all agents are activated by default. The prompt combines the template from Section 3.2 with a curated list of clinically significant findings, following prior work (tu2024towards; peng2025scaling); details appear in Appendix D.
@@
Figure 3: VQA results regarding progression.
3.5 Effectiveness of RadAgents Design
We evaluate the multi-agent design against a monolithic baseline in which a single LLM follows the ABCDE scheme (GPT-4o w/ Workflow), thereby separating improvements attributable to multi-agent coordination from those due to a structured workflow. Consistent with Table 3.1 and Figure 3, (i) introducing a workflow confers substantial gains over ReAct, and (ii) the multi-agent architecture further improves performance. These effects stem from contextual isolation within the workflow, which limits long-context drift in extended reasoning chains.
In addition, we validate the context verifier and conflict‐resolution modules. On our 1,147‐case dataset, the context verifier was triggered in 37.67% of instances. We observed a tool‐conflict rate of 32.78%, of which 78.99% were resolved correctly.
@@ 4 Discussion
In this work, we propose the first radiologist-like agentic system, which demonstrates superior performance in CXR interpretation. Further improvements are required for broader application: (1) the context verifier can only detect errors but cannot modify outputs, especially for visual evidence such as bounding boxes, so stronger supporting models are needed; (2) current experiments are limited to frontal-view images, as most existing tools are incompatible with lateral views; (3) prompting strategies could be further optimized; (4) the framework could be extended to additional modalities, such as CT and MRI; and (5) our current implementation relies on large proprietary models and multiple tools, making it costly and difficult to reproduce; exploring smaller, open-source models is a valuable direction.
@@ Appendix B Clinical Evaluation Metrics
Evaluating the quality of generated radiology reports is non-trivial. Early works adopted general-domain natural language processing metrics such as ROUGE (lin2004rouge) and BLEU (papineni2002bleu). While these metrics are widely used for text evaluation, they treat differences in wording the same as clinically significant errors, failing to reflect medical accuracy. To address this limitation, clinically informed evaluation metrics, such as CheXbert (smit2020combining), RadGraph (jain1radgraph), GREEN (ostmeier2024green), and RaTEScore (zhao2024ratescore), have been proposed to better assess clinical correctness and utility. CheXbert is based on multi-label classification results for 5 or 13 diseases (along with one extra “normal” label). RadGraph considers literal entity agreement considering the positive or negative context of each entity. GREEN judges recall and precision errors by LLM prompting. RaTEScore is inspired by RadGraph but less sensitive to phrasing by an F1-like computation which allows semantic matching between entities based on a cosine similarity. The metrics are computed using their official and standardized implementations: RadGraph-F1111https://pypi.org/project/radgraph/0.1.2/, CheXbert-F1222https://pypi.org/project/f1chexbert/, RaTE Score [^3] https://pypi.org/project/RaTEScore/0.5.0/, and GREEN [^4] https://pypi.org/project/green-score/0.0.8/.
@@ Appendix F Details of V-RAG
We retrieve images and their textual descriptions that align with the features of target medical images following (chu2025reducing). These references, rich in visual and textual medical details, guide response generation. To obtain embeddings, we use Rad-DINO, which provides robust representations across diverse CXR image types. For each image $X_{img}$, we extract its embedding $E_{img}=\mathbf{R}^{d}$, with $d=768$, and store them in the embedding memory $\mathcal{M}$.
For efficient retrieval during inference, we build $\mathcal{M}$ using FAISS  [^5] https://github.com/facebookresearch/faiss, a GPU-accelerated vector search system. We employ approximate kNN with the Hierarchical Navigable Small World (HNSW) algorithm (malkov2018efficient), enabling retrieval of the top-$k$ most similar images in $\mathcal{M}$.
@@
We study how retrieval quality changes with the number of retrieved studies $k$ with the sampled 100 cases from the MS-CXR test set used in this study. For each setting, we compute two metrics: helpful-rate, the percentage of retrieved studies that improve the answer, and harmful-rate, the percentage that hurt the answer. As shown in Table 3, As k increases, the harmful rate grows more quickly than the helpful, e.g., the helpful-rate increases from 0.48 at $k=1$ to 0.65 at $k=5$, while the harmful-rate also rises from 0.09 to 0.20. We hypothesize this is due to longer contexts imposing a heavier reasoning burden and increasing hallucination, consistent with prior LLM findings. This illustrates a trade-off: retrieving more studies provides greater chances of including helpful evidence but also increases the risk of introducing misleading content. To balance these effects, we choose $k=3$ by default, achieving a helpful-rate of 0.62 with a moderate harmful-rate of 0.14.